OpenStudy (rainbow_dashie):

Kate is making hamburgers for her friends and has 12 1/2 pounds of hamburger. How many - 1/2 pound hamburgers can she make?

OpenStudy (rainbow_dashie):

@CO_oLBoY

OpenStudy (anonymous):

- 1/2 pound hamburgers or 1/2 pound hamburgers???

OpenStudy (rainbow_dashie):

1/2

OpenStudy (anonymous):

hamburgers she can make= 12.5 / 1.5 =8.33 =8

OpenStudy (rainbow_dashie):

there is no 8

OpenStudy (anonymous):

what r the choices??

OpenStudy (rainbow_dashie):

A. 50 B. 12 1/2 C. 25 D. 2

OpenStudy (anonymous):

oh sorry its 25 coz 12.5 / 0.5 = 25

OpenStudy (rainbow_dashie):

The juice dispenser in the cafeteria holds 21 1/2 pints of juice. The glasses in the cafeteria only hold 1/4 of a pint of juice. How many full glasses can be served at breakfast?

OpenStudy (anonymous):

86

OpenStudy (rainbow_dashie):

Fredo has a coupon for $1.00 off the price of a loaf of bread at the grocery store. After he arrived at the store, he found out the bread had already been marked down $2.00. What is the total discount on the price of the bread?

OpenStudy (rainbow_dashie):

@CO_oLBoY

OpenStudy (anonymous):

$3.00

OpenStudy (rainbow_dashie):

Its wrong

OpenStudy (rainbow_dashie):

It wuz -3.00

OpenStudy (anonymous):

wait